1. Find the general solution of the differential equation using the method

of variation of parameters y′′ + 4y = 4sec(2x)



1
Expert's answer
2022-05-30T16:25:33-0400

The corresponding homogeneous differential equation is


"y''+4y=0"

Corresponding (auxiliary) equation is


"r^2+4=0"

"r_1=2i, r_2=-2i"

The general solution of the homogeneous differential equation is


"y_h=C_1\cos(2x)+C_2\sin(2x)"

Consider


"y_1=C_1\cos(2x)+C_2\sin(2x)"

"y_1'=C_1'\cos(2x)-2C_1\sin(2x)+"

"+C_2'\sin(2x)+2C_2\cos(2x)"

Let "C_1'\cos(2x)+C_2'\sin(2x)=0." Then



"y_1'=-2C_1\sin(2x)+2C_2\cos(2x)"

"y_1''=-2C_1'\sin(2x)-4C_1\cos(2x)"

"+2C_2'\cos(2x)-4C_2\sin(2x)"

Substitute


"-2C_1'\sin(2x)-4C_1\cos(2x)"

"+2C_2'\cos(2x)-4C_2\sin(2x)"




"+4C_1\cos(2x)+4C_2\sin(2x)=4\sec (2x)"

Solve the system


"C_1'\cos(2x)+C_2'\sin(2x)=0""-2C_1'\sin(2x)+2C_2'\cos(2x)=\dfrac{4}{\cos(2x)}"

"C_1'\cos(2x)+C_2'\sin(2x)=0""-C_1'\sin(2x)\cos(2x)+C_2'\cos^2(2x)=2"

"C_1'\cos(2x)+C_2'\sin(2x)=0""C_2'=2"


"C_1'=-2\dfrac{\sin(2x)}{\cos(2x)}""C_2'=2"

Integrate


"C_1=-2\int \dfrac{\sin(2x)}{\cos(2x)}dx=\ln|\cos(2x)|+C_3"

"C_2=2x+C_4"

The general solution of the non homogeneous differential equation is


"y=\ln|\cos(2x)|\cos(2x)+C_3\cos(2x)""+2x\sin(2x)+C_4\sin(2x)"